OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line parallel to y = –8x – 3 through (–1, 4)

OpenStudy (anonymous):

y-4=-8(x+1) y-4=-8x-8 y=-8x-8+4 y=-8x-4